OpenStudy (anonymous):

Solve the inequality. 2x+1/3 <9

OpenStudy (anonymous):

The answer is 5 2/5 right?

OpenStudy (anonymous):

Or can someone correct me ?

OpenStudy (anonymous):

\[2x + \frac{ 1 }{ 3 } < 9\]

OpenStudy (muzzack):

2x<9-1/3

OpenStudy (anonymous):

Umm soo?

OpenStudy (anonymous):

Get x on one side of the equation by itself.

OpenStudy (anonymous):

I know i did that and my answer was 5 2/5 ...am i right?

OpenStudy (anonymous):

\[2x < 9 - \frac{ 1 }{ 3 }\] \[2x < \frac{ 26 }{ 3 }\] \[x < \frac{ 13 }{ 3 }\]

OpenStudy (anonymous):

Ohh okay so 4 1/3?

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

Thanks ! :)